The Number of Monomer-Dimer Tilings of an m by n Rectangle with a Fixed Proportion of Monomers

Gleb Pogudin and Doron ZEILBERGER

Abstract:

Supporting Maple Package and Output All the results in this article were obtained by the use of the Maple packages

• http://www.math.rutgers.edu/~zeilberg/tokhniot/BDD.txt ,

whose output files, along with links to diagrams, are available from the front of this article

http://www.math.rutgers.edu/~zeilberg/mamarim/mamarimhtml/bdd.html .

Acknowledgment References

Gleb Pogudin, RISC-Linz Email: .

Doron Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA.

Email: DoronZeil at gmail dot com .